OpenStudy (anonymous):

the product of -5/9 and -1/13 divided by 1/11, translate to number

OpenStudy (anonymous):

\[\frac{ -5 }{ 9 }*\frac{ -1 }{ 13 }*\frac{ 11 }{ 1 } = \frac{ 55 }{ 117 }\]

OpenStudy (anonymous):

\[\frac{\frac{-5}{9}\times \frac{-1}{13}}{\frac{1}{11}}=\frac{\frac{-5\times -1}{9\times 13}}{\frac{1}{11}}\]\[=\frac{-5\times -1}{9\times 13}\times \frac{11}{1}\]